Solve Probability Problems. P(x>450) Mean=390 Deviation=65
Answer by stanbon(75887)

Solve Probability Problems. P(x>450) Mean=390 Deviation=65
-----
z(450) = (450-390)/65 = 60/65 = 12/13
---
P(x > 45) = P(z > 12/13) = normalcdf(12/13,100) = 0.1780
=======================
Cheers,
Stan H.
===============